OpenStudy (anonymous):

Write the expression ln (2/ sqrt(x)) - ln sqrt(2x) as a single logarithm

OpenStudy (anonymous):

\[A. \ln(\sqrt{2}/x) B. \ln (2^{3/2}\sqrt{x}) C. \ln(\sqrt{2x}) D. \ln (x/\sqrt{2}) \] \[E. \ln (x ^{3/2}\sqrt{2})\]

OpenStudy (anonymous):

\[\ln (2/\sqrt{x}/\sqrt{2x})\]

OpenStudy (anonymous):

so its A

OpenStudy (anonymous):

\[\ln (2/\sqrt{2}x)\]

OpenStudy (anonymous):

\[\ln (\sqrt{2}/x)\]